OpenStudy (anonymous):

can anyone solve these problems??? M,A,M,J,J,A,S,O,..........What comes next? a.N b.M c.A d.P

OpenStudy (anonymous):

sathi, you posted a wrong problem, whats wrong with you,

OpenStudy (anonymous):

the sequence is 3, 7, 31, 211

OpenStudy (anonymous):

NOT 3, 7, 32, 211,

OpenStudy (anonymous):

so how did u prove that 3 , 7 31, 211

OpenStudy (anonymous):

but the question was same in the exam paper. 31 not 32

OpenStudy (anonymous):

but anyway how did u solve that ..

OpenStudy (anonymous):

you posted 32 , scroll to your question

OpenStudy (anonymous):

look at the problem you posted earlier you will find 32

OpenStudy (anonymous):

oh... i mean 32 not 31

OpenStudy (anonymous):

whats ur process

OpenStudy (anonymous):

no, wait, admit that you posted an incorrect problem

OpenStudy (anonymous):

ok...

OpenStudy (anonymous):

look at the post , on the left, scroll down

OpenStudy (anonymous):

follow me , one sec

OpenStudy (anonymous):

wow i got answer

OpenStudy (anonymous):

ok so the sequence 3, 7, 31, 211, is 1 + 2, 1 + 2 * 3 , 1 + 2 * 3 * 5 , 1 + 2*3*5*7

OpenStudy (anonymous):

its option no.A...N

OpenStudy (anonymous):

wait, lets do the sequence

OpenStudy (anonymous):

what is special about 2,3,5,7

OpenStudy (anonymous):

they are primes

OpenStudy (anonymous):

so the next number will be 1 + 2*3*5*7*11 = 2311

OpenStudy (anonymous):

2311, 30031, 510511

OpenStudy (anonymous):

yea i agree with ur idea...thanks

OpenStudy (anonymous):

how did you solve your MAMJAN problem?

OpenStudy (anonymous):

its the series of month starting from M(march)

OpenStudy (anonymous):

is that a code, letters mean numbers

OpenStudy (anonymous):

haha

OpenStudy (anonymous):

nice one

OpenStudy (anonymous):

X % of Y is Y % of : a.x b.y/100 c.x/100 d.100x

OpenStudy (anonymous):

wait, this is an open question, the sequence before are called euclid numbers

OpenStudy (anonymous):

are there an an infinite number of prime euclid numbers, E6 is not prime for example

OpenStudy (anonymous):

hey man m the student of medicine so i dont know the mathmatical calculation much more......so would u plz describe in simple way.

OpenStudy (anonymous):

my guess was a x

OpenStudy (anonymous):

yes a

OpenStudy (anonymous):

x is correct

OpenStudy (anonymous):

but its c in book

OpenStudy (anonymous):

x% of y = x/ 100 * y = x * y/100 = y% of x

OpenStudy (anonymous):

oh , hmmmm

OpenStudy (anonymous):

well another question.

OpenStudy (anonymous):

in a class with a certain number of students , if one student weighing 50kg is added then average weight of class incereases by 1 kg ,if one more student weighin 50kg is added then the average weight of class increasses by 1.5kg over the original average? what isthe original average weight (in kg) of the class?